OpenStudy (anonymous):

solve: 27^(2x+1)=9^(4x)

OpenStudy (anonymous):

o i think i fanned you there was a blue arrrow next to ur name i clicked it

OpenStudy (anonymous):

Oh lol nice

OpenStudy (anonymous):

3^(3*(2x+1))=3^(2*4x) 3*(2x+1)=2*4x 6x+3=8x x=3/2